OpenStudy (anonymous):

Which of the following is the slope between the points (-6, 0) and (2, -9)? eight ninths -eight ninths nine eighths -nine eighths

OpenStudy (anonymous):

point slope formula

OpenStudy (johnweldon1993):

for this...you use the formula...(y2-y1)/(x2-x1)

OpenStudy (anonymous):

\[\frac{ Y_2-Y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

haha good job everyone!

OpenStudy (johnweldon1993):

haha at least she know's everyones right because they're on the same page

OpenStudy (anonymous):

indeed

OpenStudy (johnweldon1993):

And @tristynhart remember coordinates are in form (x,y)

OpenStudy (anonymous):

@tristynhart did you get it? rather, do you understand it?

OpenStudy (johnweldon1993):

@tristynhart need a little more?

OpenStudy (johnweldon1993):

@e.cociuba no the answer is there...I just don't know if she needs a bit clearer of an explaination

OpenStudy (johnweldon1993):

Alright last hint will be (-6, 0) and (2, -9) x1 y1 x2 y2

OpenStudy (anonymous):

(-6, 0) and (2, -9)? y2 = -9 y1 = 0 x2 = 2 x1 = -6 (y2-y1)/(x2-x1) = ??